OpenStudy (anonymous):

A prism has a right triangle as its base with a hypotenuse that measures 5 meters and legs that measure 3 meters and 4 meters. What is the surface area of this prism if the height of the prism is 6 meters? Answer 78 m2 84 m2 96 m2 102 m2

OpenStudy (hoblos):

lol again... it is 84

OpenStudy (anonymous):

Perimeter of the base \( \times \) height \(+ 2 \times\) area of the base

OpenStudy (anonymous):

so hoblos gave you the right answer :)

OpenStudy (anonymous):

thank u guys again

OpenStudy (anonymous):

Glad to help again (sunset)

OpenStudy (anonymous):

\[ (5+3+4) \times 6 + 2 \times (\frac 12 \times 3 \times 4) =84\]

OpenStudy (anonymous):

i have another one this one is for a pyramid ill post it if u want to help

OpenStudy (anonymous):

New thread :)